OpenStudy (anonymous):

Is this standard form for the parabola x^2=y+8x=16?

OpenStudy (anonymous):

\[x ^{2}=y+8x-16\] \[(x-4)^{2}=(y-16)\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

or\[(x-0)^{2}=4(2)(y-16)\]

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@dan815 @aajugdar

OpenStudy (sweetburger):

y=ax^2+bx+c is standard form right?

OpenStudy (anonymous):

for a parabola standard form is (y-k)^2=4p(x-h) or (x-h)^2=4p(y-k)

OpenStudy (kropot72):

The general equation of a parabola is: \[y ^{2}=4ax\ ...........(1) \] The equation that you posted (y-k)^2=4p(x-h) is the curve (1) given a translation of \[\left(\begin{matrix}h \\ k\end{matrix}\right)\]

OpenStudy (anonymous):

but thats what the textbook wants me to use to create an equation-standard form equation of the parabola

OpenStudy (kropot72):

Please post the exact question from the textbook.

OpenStudy (anonymous):

Identify the conic section represented by each equation. Write the equation in standard form and graph the equation. \[x ^{2}=y+8x-16\]

OpenStudy (kropot72):

If the curve of the parabola \[y ^{2}=4ax\] is reflected in the line y = x we get the graph of the inverse curve which is: \[y=\frac{1}{4a}x ^{2}\]

OpenStudy (kropot72):

The given equation is \[x ^{2}=y+8x-16\] Rearranging gives \[y=x ^{2}-8x+16\] which factorizes to \[y=(x-4)^{2}\ .............(2)\] (2) is the equation of a parabola with vertex (4, 0 and translation \[\left(\begin{matrix}4 \\ 0\end{matrix}\right)\] When x = 0, y = 16 and when y = 0, x = 4. The sketch of the curve follows: